On two examinations, you have grades of 89 and 83. there is an optional final examination, which counts as one grade. you decide to take the final in order to get a course grade of a, meaning a final average of at least 90. What must you get on the final to earn an A in the course? By taking the final, if you do poorly, you might risk the B that you have in the course based on the first two exam grades. If your final average is less than 80, you will lose your B in the course. Describe the grades on the final that will cause this to happen.
Answer by macston(5194)

For the average of 3 grades to be an A ( 90 or above) the sum of the grades must be at least 3(90)=270. Current grades are 89+83=172 and 270-172=98, so your grade on the final must be 98 or better for a minimum 90 average. However if the average of the 3 grades falls below 80 (the sum of the grades <3(80) or 240) then you will receive a C. the sum of your current grades is 172 So since 240-172=68 if your grade on the final is below 68 points, you will receive a C.
SUMMARY you must score 98 or better on the final to receive an A in the course,
a score of 68 to 98 retains the B, but a score below 68 will lower your grade to a C.